OpenStudy (anonymous):

A right triangle has base (x - 6) units and height (6x - 12) units. Part A: What is the square of the length of the hypotenuse of the triangle? Show your work. (4 points) Part B: What is the area of the triangle? (3 points) Part C: Using the solution obtained in Part B, explain the closure property of multiplication of polynomials. (3 points)

OpenStudy (anonymous):

that would be \[(x-6)^2+(6x-12)^2\] for the square of the length of the hypotenuse, since via pythagoras \[a^2+b^2=c^2\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What is B

OpenStudy (jdoe0001):

\(\bf \textit{area of a triangle}=\cfrac{1}{2}\cdot base\cdot height\)

OpenStudy (anonymous):

I dont know the answer

OpenStudy (jdoe0001):

well, you have the base, and the height, so, is all you need

OpenStudy (jdoe0001):

"A right triangle has base \(\bf (x - 6)\) units and height \(\bf (6x - 12)\) units."

OpenStudy (anonymous):

ok